OpenStudy (anonymous):

so...6u+1=0 u= 1 over -6 and 3u-1=0 so u= 1 over -3? Right......

OpenStudy (anonymous):

6u+1=0 6u=-1 u=-1/6 3u-1=0 3u=1 u=1/3

OpenStudy (anonymous):

aka correct.

OpenStudy (anonymous):

LOL Thanks ^^

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Sorry, your second answer was off. Fix that, THEN both are correct :P

OpenStudy (anonymous):

Oke XD